OpenStudy (anonymous):

If two cubes have edges of 1 and 2, what is the sum of their volumes??

OpenStudy (anonymous):

x*x*x=volume

OpenStudy (anonymous):

the volume of a cube is its length x width x height... for a cube... it's (edge)^3

OpenStudy (anonymous):

first cube x=1 second cube x=2

OpenStudy (anonymous):

x is one of the sides

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

volume1 = 1*1*1 volume2 = 2*2*2

OpenStudy (anonymous):

So it is 9

OpenStudy (anonymous):

Multiply it out and add them together.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright thank you!